Last bits of set-up continue, PI's laser and CTEMPS fiber are still out of commission, so there isn't an urgency for us to finish.
Today:
- Installed all sensors on aux towers. One CSAT head (S/N 2048) didn't turn "green", so replaced with 1007 (uconv.a1).
 - Began to clean up the aux tower cable mess, but a quite a bit more to go.
 - Added an extension to the 9.3m CSAT cable at lconv. (You can relax now, Gary.)
 - Installed soil sensors and rad tripod in the soybeans at lconv, hooked up to lconv.a2, port 7. TRH booms didn't fit the tripod U-bolt mount, so purchased smaller stock tonight in town.
 - My intent now is to install a rad tripod in the grass at uconv (the most exposure for the radiometers to grass) and install a 10' rohn in the corn at rel (into rel.a1 port 7).
 - Plugged in power monitor to lconvt. Will have to move a battery box at rel to connect the usb cable to pmon there.
 - Created a new CSAT box tower in the storage locker.
 - Had a great lunch at the Chinese place beside the Mahomet Ace hardware.
 
Plan for tomorrow:
- add radiometer and boom to lconv beans
 - install rad/soils in grass at uconv
 - install rad/soils in corn at rel
 - maybe move lconv 6m particle counter to other tower face, though isn't terrible where it is.
 - Organize the base trailer (though getting better already, as we've moved things out and Gary hung the monitor).
 - start looking at data! (I expect to see <some> problems).